175 mL of 0.25 M NaNO3
1
Expert's answer
2016-11-25T08:35:10-0500
Molarity = quantity of moles divided by liters.
M = 0.25M
V = 175 mL = 0.175L
n = 0.175*0.25 = 0.04375 mole
mass = quantity of moles multiplied by molar mass;
m = 0.04375*85=3.17875 grams
Answer: 3.17875 g; 0.04375 moles.